# Lexicon3D: Probing Visual Foundation Models for Complex 3D Scene Understanding

### v1

@@ Abstract
Complex 3D scene understanding has gained increasing attention, with scene encoding strategies playing a crucial role in this success. However, the optimal scene encoding strategies for various scenarios remain unclear, particularly compared to their image-based counterparts. To address this issue, we present a comprehensive study that probes various visual encoding models for 3D scene understanding, identifying the strengths and limitations of each model across different scenarios. Our evaluation spans seven vision foundation encoders, including image-based, video-based, and 3D foundation models. We evaluate these models in four tasks: Vision-Language Scene Reasoning, Visual Grounding, Segmentation, and Registration, each focusing on different aspects of scene understanding. Our evaluations yield key findings: DINOv2 demonstrates superior performance, video models excel in object-level tasks, diffusion models benefit geometric tasks, and language-pretrained models show unexpected limitations in language-related tasks. These insights challenge some conventional understandings, provide novel perspectives on leveraging visual foundation models, and highlight the need for more flexible encoder selection in future vision-language and scene-understanding tasks.
@@ 1 Introduction
Recently, complex 3D scene understanding has emerged as a pivotal area in computer vision, encompassing tasks such as scene generation [74, 33, 25, 24, 26], reasoning [35, 55, 5, 52], and interaction [36, 108]. Leveraging large-scale vision foundation models, approaches like [91, 64, 68, 84, 42] have achieved promising results, thereby enabling a wide range of real-world applications, from autonomous driving [54, 75, 79, 112], robotics [108, 57], to multi-modal agents [78, 1].
While numerous studies [6, 99, 67] have provided guidance on the use of vision foundation models for 2D image-based tasks, the strategies for 3D scenarios remain unclear. A systematic understanding of complex real-world scenarios involves not only semantic and depth awareness [6], which is possible to evaluate within the 2D domain, but also geometric awareness and the ability to align with multi-modal information for reasoning and grounding tasks. To address this gap, our work evaluates the use of different types of visual foundation models for complex scene understanding and seeks to identify the strengths and limitations of each model in different scenarios. Ultimately, this study aims to contribute to the development of more effective and efficient scene understanding systems.
@@
To answer these questions, we design a unified paradigm to systematically probe visual encoding models for complex 3D scene understanding from different perspectives. Our evaluation spans seven vision foundation models in images, videos, and 3D-based ones, as shown in Table 1. Our evaluations are conducted among four diverse tasks: Vision-Language Scene Reasoning assesses the model’s ability to reason about scenes based on textual descriptions, evaluating scene-level representation; Visual Grounding tests the model’s capacity to associate language with specific objects within a scene, reflecting object-level representation; Segmentation evaluates the model’s ability to assign semantic labels to each pixel, assessing semantic understanding; Registration measures the performance of aligning different views of a scene, testing geometric capacity. Through these tasks, our aim is to explore the strengths and weaknesses of different vision foundation models in 3D scene understanding, providing insights into their applicability in various scenarios. With the major results demonstrated in Figure 1, our key findings include:
  • Image or video foundation models achieve promising results for 3D scene understanding. Among them, DINOv2 [58] demonstrates the best overall performance, showing strong generalizability and flexibility, which is consistent with the observation in 2D [6]. Our evaluation further verifies its capability in global and object-level 3D vision-language tasks. It can serve as a general backbone for 3D scene understanding.
@@
  • Visual encoders pretrained with language guidance do not necessarily perform well in other language-related evaluation tasks, challenging the common practice of using such models as default encoders for vision-language reasoning tasks.
@@
We name our work as Lexicon3D, a unified probing architecture and the first comprehensive evaluation of 3D scene understanding with visual foundation models. The key findings we have achieved above, in conjunction with other interesting observations, suggest exploring more flexible encoder selections in future vision-language tasks to optimize performance and generalization.
Figure 1: Evaluation settings and major results of different vision foundation models for complex 3D scene understanding. We probe visual foundation models of different input modalities and pretraining objectives, assessing their performance on multi-modal scene reasoning, grounding, segmentation, and registration tasks.
@@
Swin3D [93] 3D Points Swin3D-L Annotation Structure3D
Table 1: Details of the seven evaluated visual encoding models.
@@ 2 Related Work
Our work is closely related to methods that focus on extraction of features from images, videos, and 3D assets, as well as learning joint spaces for vision-language fusion. A large body of recent literature has explored the representation learning for multi-modal visual inputs and their complementary performance in image understanding. In contrast, our paper presents a comprehensive analysis of the use of pretrained visual encoders for zero-shot 3D scene understanding. To the best of our knowledge, we are the first to examine pretrained video encoders on 3D scene understanding tasks and to compare image, video, and 3D point encoding strategies in this context.
@@
Image self-supervised learning.   In recent years, learning robust and generalizable pretrained image representations has become a prevalent research direction in computer vision and multi-modal research. One line of work focuses on learning task-agnostic image features using self-supervised learning (SSL) signals, which include pretext tasks such as colorization [100], inpainting [62], transformation prediction [27], and self-distillation [30, 19, 18, 29, 14]. The recent development of the patch-based image tokenizer, ViT [22], has also led to the emergence of mask autoencoder architectures (MAE) for feature extraction [8, 31, 111]. Of particular interest, DINOv2 [58], combining a masked-image modeling loss and an invariance-based self-distillation loss, has become one of the most scalable and competitive self-supervised learning architectures using only image signals. Another line of work proposes learning image features with text guidance, i.e., using textual descriptions to guide the pretraining of the image encoders [38, 53]. Building upon the powerful image-text encoder CLIP [65], LSeg [44] and BLIP [45, 46] extend the image pretraining objective to more complex visual perception tasks by incorporating pixel-level semantic understanding and encouraging better alignment with large language models (LLMs) [13, 103, 66, 102], respectively.
@@
Video and 3D representation learning.   Self-supervised representation learning has also been explored in the context of videos and 3D point clouds. Extending the success of the CLIP architecture [65] from images to videos, a body of work proposes to pretrain a video encoder by aligning the feature space with textual guidance extracted from video captions [89, 97, 3, 85]. Other pretext tasks used in video representation learning include next frame prediction [10] and MAE [80, 83, 28]. Among them, Bardes et al. [11] adapt the MAE-inspired joint embedding prediction architecture (JEPA) [4, 43] to the spatio-temporal domain, achieving state-of-the-art performance on a wide spectrum of video and image tasks. Despite the extensive research on 2D visual foundation encoders, pretrained models for 3D point clouds are significantly fewer due to the lack of large-scale 3D datasets. Existing work has explored contrastive pretraining [37, 88, 105] and masked signal modeling [48, 59, 87, 92, 96, 101] for point representation learning. Recently, benefiting from the rapid advancement of 3D data rendering and a large synthetic dataset [109], Swin3D [93] has outperformed other pretraining methods by a significant margin using supervised pretraining.
@@ 3 Probing Visual Encoders for Scene Understanding
The objective of our Lexicon3D is to evaluate different visual foundation models in complex scene understanding tasks. We first construct a unified architecture capable of probing different visual foundation models on a spectrum of downstream tasks. Then, we break down the 3D scene understanding task into four sub-tasks, including (1) vision-language reasoning, (2) visual grounding, (3) semantic understanding, and (4) geometric understanding, for a more detailed evaluation.
@@ 3.1 A Unified Probing Framework
Figure 2: Our unified probing framework to evaluate visual encoding models on various tasks.
Figure 3: Visualizations of extracted scene features from different visual encoders using PCA. The clear distinction between colors and patterns demonstrates the behaviors of different models.
We design a unified framework, as shown in Figure 2, to extract features from different foundation models, construct a 3D feature embedding as scene embeddings, and evaluate them on multiple downstream tasks. For a complex indoor scene, existing work usually represents it with a combination of 2D and 3D modalities. For realistic scenarios [20, 94, 15], videos are usually first captured with handheld cameras and then 3D points are obtained from reconstruction algorithms such as COLMAP [72]. For digital and synthetic scenarios [69, 109], 3D assets are designed and generated first, before images and/or videos are rendered within the created space. Given a complex scene represented in posed images, videos, and 3D point clouds, we extract their feature embeddings with a collection of vision foundation models. For image and video-based models, we project their features into the 3D space for subsequent 3D scene evaluation tasks with a multi-view 3D projection module. Following [35, 63, 34, 21], for a point cloud $\mathbf{P}$, this module produces features $f_{\mathbf{p}}$ for each point $\mathbf{p}\in\mathbf{P}$ given image features $f$ and the pose and camera information $\mathbf{K,R}$. We first project all points onto the image plane to obtain their corresponding pixel features. Concretely, for a point $\mathbf{p}$, we obtain its projected pixel $\mathbf{u}$ on the image $i$ with
$\tilde{\mathbf{u}}=\mathbf{K}_{i}\mathbf{R}_{i}\tilde{\mathbf{p}},\quad\tilde{\mathbf{u}},\tilde{\mathbf{p}}\text{\ represent homogeneous coordinates of }\mathbf{u},\mathbf{p}.$ (1)
@@
After projection, we obtain 3D feature fields represented as point cloud feature embeddings for each visual foundation model. We use these as input to the shallow probing heads to evaluate various downstream tasks. To minimize the effect of the model finetuning process, we freeze the parameters for the encoding models to be evaluated, and only tune the linear or shallow probing head for all tasks.
@@
Models.   In this work, we primarily focus on evaluating visual foundation models that are frequently leveraged by recent complex scene understanding and multi-modal reasoning models. A complex scene can often be represented in posed 2D images and videos or in 3D point clouds. The image and video modalities sacrifice explicit geometry information, but they preserve rich and dense semantic and textural information of a scene. Conversely, the point cloud modality offers the opposite trade-offs. Additionally, the 2D modalities benefit from strong foundation models trained on vast amounts of data, while 3D point backbones only leverage much smaller datasets.
@@
We categorize visual foundation models into three categories, with an overview of the evaluated models provided in Table 1. For image encoders, we evaluated DINOv2 [58], LSeg [44], CLIP [65], and StableDiffusion (SD) [70]. For the video modality, we evaluate V-JEPA [11], the state-of-the-art video understanding model succeeding VideoMAE [80, 83] for a wide spectrum of perception and reasoning tasks, as well as StableVideoDiffusion (SVD) [12], a video generative model. The lack of large-scale 3D scene-level datasets hinders the development of strong zero-shot generalizable 3D foundation models as opposed to their 2D counterparts. However, for comparison, we evaluated Swin3D [93], a 3D backbone that achieves leading performance in zero-shot perception tasks in multiple evaluation datasets compared to earlier methods [88, 37, 105]. Swin3D is pretrained on Structured3D [109], a dataset 10 times larger than ScanNet [20].
@@
Feature visualization.   Figure 3 visualizes the scene features extracted by the vision foundation models. To visualize a high-dimensional feature space with $C$ channels, we apply principal component analysis (PCA) to reduce the feature dimensions to three, normalize them to range $[0,1]$, and interpret them as RGB color channels. The visualizations reveal several intuitive findings. The image models, DINOv2 and LSeg, demonstrate strong semantic understanding, with LSeg exhibiting clearer discrimination due to its pixel-level language semantic guidance. The diffusion-based models, SD and SVD, in addition to their semantic modeling, excel at preserving the local geometry and textures of the scenes, because of the generation-guided pretraining. The video models, SVD and V-JEPA, showcase a unique ability to identify different instances of the same semantic concepts, such as the two trees in the first scene and the chairs in both scenes. The 3D model, Swin3D, also exhibits strong semantic understanding. However, due to limited training data and domain shift, its quality is not on par with the image foundation models, despite being pretrained on perfect semantic annotations.

### v2

@@ Abstract
Complex 3D scene understanding has gained increasing attention, with scene encoding strategies built on top of visual foundation models playing a crucial role in this success. However, the optimal scene encoding strategies for various scenarios remain unclear, particularly compared to their image-based counterparts. To address this issue, we present the first comprehensive study that probes various visual encoding models for 3D scene understanding, identifying the strengths and limitations of each model across different scenarios. Our evaluation spans seven vision foundation encoders, including image, video, and 3D foundation models. We evaluate these models in four tasks: Vision-Language Scene Reasoning, Visual Grounding, Segmentation, and Registration, each focusing on different aspects of scene understanding. Our evaluation yields key intriguing findings: Unsupervised image foundation models demonstrate superior overall performance, video models excel in object-level tasks, diffusion models benefit geometric tasks, language-pretrained models show unexpected limitations in language-related tasks, and the mixture-of-vision-expert (MoVE) strategy leads to consistent performance improvement. These insights challenge some conventional understandings, provide novel perspectives on leveraging visual foundation models, and highlight the need for more flexible encoder selection in future vision-language and scene understanding tasks.
@@ 1 Introduction
Recently, complex 3D scene understanding has emerged as a pivotal area in computer vision, encompassing tasks such as scene generation [77, 34, 26, 25, 27, 96], reasoning [36, 58, 5, 55], and interaction [37, 112]. Leveraging large-scale vision foundation models, many approaches [94, 67, 71, 87, 44] have achieved promising results in various downstream tasks, thereby enabling a wide range of real-world applications, from autonomous driving [57, 78, 82, 117], robotics [112, 60], to multimodal agents [81, 1].
While numerous studies [6, 103, 70] have provided guidance on the use of vision foundation models for 2D image-based tasks, the strategies for 3D scenarios remain unclear. A systematic understanding of complex real-world scenarios involves not only semantic and depth awareness [6], which is possible to evaluate within the 2D domain, but also geometric awareness and the ability to align with multimodal information for reasoning and grounding tasks. To address this gap, our work evaluates the use of different types of visual foundation models for complex scene understanding and seeks to identify the strengths and limitations of each model in different scenarios. Ultimately, this study aims to contribute to the development of more effective and efficient scene understanding systems.
@@
To answer these questions, we design a unified paradigm to systematically probe visual encoding models for complex 3D scene understanding from different perspectives. Our evaluation spans seven vision foundation models in images, videos, and 3D-based models, as shown in Table 1. Our evaluation is conducted among four diverse tasks: Vision-Language Scene Reasoning assesses the model’s ability to reason about scenes based on textual descriptions, evaluating scene-level representation; Visual Grounding tests the model’s capacity to associate language with specific objects within a scene, reflecting object-level representation; Segmentation evaluates the model’s ability to assign semantic labels to each pixel, assessing semantic understanding; Registration measures the performance of aligning different views of a scene, testing geometric capacity. Through these tasks, our aim is to explore the strengths and weaknesses of different vision foundation models in 3D scene understanding, providing insights into their applicability in various scenarios. With the major results demonstrated in Figure 1, our key findings include:
  • Image or video foundation models achieve promising results for 3D scene understanding. Among them, DINOv2 [61] demonstrates the best overall performance, showing strong generalizability and flexibility, which is consistent with the observation in 2D scenarios [6]. Our evaluation further verifies its capability in global and object-level 3D vision-language tasks. It can serve as a general backbone for 3D scene understanding.
@@
  • Visual encoders pretrained with language guidance (e.g., CLIP [68]) do not necessarily perform well in other language-related evaluation tasks, challenging the common practice of using such models as default encoders for vision-language reasoning tasks.
@@
  • The mixture-of-vision-expert (MoVE) strategies, including combining multi-layer features from the same visual model, and concatenating features from multiple visual models, both lead to a consistent boost of performance across different tasks.
Our work, Lexicon3D, provides a unified probing architecture and the first comprehensive evaluation of 3D scene understanding with visual foundation models. The key findings we have achieved above, in conjunction with other interesting observations, suggest exploring more flexible encoder selections in future vision-language tasks to optimize performance and generalization.
Figure 1: Evaluation settings and major results of different vision foundation models (VFMs) for complex 3D scene understanding. We assess the performance of VFMs on multimodal scene reasoning, grounding, segmentation, and registration tasks.
@@
Swin3D [97] 3D Points Swin3D-L Segmentation Structure3D
Table 1: Details of the seven evaluated VFMs. In supervision signals, we use “SSL” to represent self-supervised learning, and use “VLM” to represent vision-language modality alignment. A more detailed explanation of the evaluated VFMs is provided in the supplementary material A.
@@ 2 Related Work
Our work is closely related to methods that focus on extraction of features from images, videos, and 3D assets, as well as learning joint representation spaces for vision-language fusion. A large body of recent literature has explored the representation learning for multimodal visual inputs and their complementary performance in image understanding. In contrast, our study presents a comprehensive analysis of the use of pretrained visual encoders for zero-shot 3D scene understanding. To the best of our knowledge, we are the first to examine pretrained video encoders on 3D scene understanding tasks and to compare image, video, and 3D point encoding strategies in this context.
@@
Image self-supervised learning.   In recent years, learning robust and generalizable pretrained image representations has become a prevalent research direction in computer vision and multimodal research. One line of work focuses on learning task-agnostic image features using self-supervised learning (SSL) signals, which include pretext tasks such as colorization [104], inpainting [65], transformation prediction [28], and self-distillation [31, 19, 18, 30, 14]. The recent development of the patch-based image tokenizer, ViT [23], has also led to the emergence of mask autoencoder architectures (MAEs) for feature extraction [8, 32, 115]. Of particular interest, DINOv2 [61], combining a masked-image modeling loss and an invariance-based self-distillation loss, has become one of the most scalable and competitive self-supervised learning architectures that uses only image signals. Another line of work proposes learning image features with text guidance, i.e., using textual descriptions to guide the pretraining of the image encoders [39, 56]. Building upon the powerful image-text encoder CLIP [68], LSeg [46] and BLIP [47, 48] extend the image pretraining objective to more complex visual perception tasks by incorporating pixel-level semantic understanding and encouraging better alignment with large language models (LLMs) [13, 107, 69, 106], respectively.
@@
Video and 3D representation learning.   Self-supervised representation learning has also been explored in the context of videos and 3D point clouds. Extending the success of the CLIP architecture [68] from images to videos, a body of work proposes to pretrain a video encoder by aligning the feature space with textual guidance extracted from video captions [92, 101, 3, 88]. Other pretext tasks used in video representation learning include next frame prediction [10] and MAE [83, 86, 29]. Among them, V-JEPA [11] adapts the MAE-inspired joint embedding prediction architecture (JEPA) [4, 45] to the spatio-temporal domain, achieving state-of-the-art performance on a wide spectrum of video and image tasks. Despite extensive research on 2D visual foundation encoders, pretrained models for 3D point clouds are significantly fewer due to the lack of large-scale 3D datasets. Existing work has explored contrastive pretraining [38, 91, 109] and masked signal modeling [50, 62, 90, 95, 100, 105] for point representation learning. Recently, benefiting from the rapid advancement of 3D data rendering and large synthetic datasets [113, 21], Swin3D [97] and Uni3D [116] have outperformed other pretraining methods by a significant margin with large-scale pretraining for scene-level perception and object-level understanding, respectively.
@@ 3 Probing Visual Encoders for Scene Understanding
The objective of Lexicon3D is to evaluate different visual foundation models in complex scene understanding tasks. We first construct a unified architecture capable of probing different visual foundation models on a spectrum of downstream tasks. Then, we break down the 3D scene understanding task into four sub-tasks, including (1) vision-language reasoning, (2) visual grounding, (3) semantic understanding, and (4) geometric understanding, for a more detailed evaluation.
@@ 3.1 A Unified Probing Framework
Figure 2: Our unified probing framework to evaluate visual foundation models on various tasks.
Figure 3: Visualization of extracted scene features from different visual foundation models. We use principal component analysis (PCA) to compress the feature embeddings into three dimensions. The clear distinction between colors and patterns demonstrates the behaviors of different models.
We design a unified framework, as shown in Figure 2, to extract features from different foundation models, construct a 3D feature embedding as scene embeddings, and evaluate them on multiple downstream tasks. For a complex indoor scene, existing work usually represents it with a combination of 2D and 3D modalities. For realistic scenarios [20, 98, 15], videos are usually first captured with handheld cameras and then 3D points are obtained from reconstruction algorithms such as COLMAP [75]. For digital and synthetic scenarios [72, 113], 3D assets are designed and generated first, before images and/or videos are rendered within the created space. Given a complex scene represented in posed images, videos, and 3D point clouds, we extract their feature embeddings with a collection of vision foundation models. For image- and video-based models, we project their features into the 3D space for subsequent 3D scene evaluation tasks with a multi-view 3D projection module. Following [36, 66, 35, 22], for a point cloud $\mathbf{P}$, this module produces features $f_{\mathbf{p}}$ for each point $\mathbf{p}\in\mathbf{P}$ given image features $f$ and the pose and camera information $\mathbf{K,R}$. We first project all points onto the image plane to obtain their corresponding pixel features. Concretely, for a point $\mathbf{p}$, we obtain its projected pixel $\mathbf{u}$ on the image $i$ with
$\tilde{\mathbf{u}}=\mathbf{K}_{i}\mathbf{R}_{i}\tilde{\mathbf{p}},\quad\tilde{\mathbf{u}},\tilde{\mathbf{p}}\text{\ represent homogeneous coordinates of }\mathbf{u},\mathbf{p},\text{respectively}.$ (1)
@@
After projection, we obtain 3D feature fields represented as point cloud feature embeddings for each VFM, and use them as input to the shallow probing heads to evaluate various downstream tasks. To minimize the effect of the model finetuning process, we freeze the parameters for the encoding models to be evaluated, and only tune the linear or shallow probing heads for all tasks.
@@
Models.   In this work, we focus primarily on evaluating visual foundation models that are frequently leveraged by recent complex scene understanding and multimodal reasoning models. A complex scene can often be represented in posed 2D images and videos or in 3D point clouds. The image and video modalities sacrifice explicit geometry information, but they preserve rich and dense semantic and textural information of a scene. Conversely, the point cloud modality offers the opposite trade-offs. Additionally, the 2D modalities benefit from strong foundation models trained on vast amounts of data, while 3D point backbones only leverage much smaller datasets.
@@
We categorize visual foundation models into three categories, with an overview of the evaluated models provided in Table 1. For image encoders, we evaluate DINOv2 [61], LSeg [46], CLIP [68], and StableDiffusion (SD) [73]. For the video modality, we evaluate V-JEPA [11], the state-of-the-art video understanding model succeeding VideoMAE [83, 86] for a wide spectrum of perception and reasoning tasks, as well as StableVideoDiffusion (SVD) [12], a video generative model. The lack of large-scale 3D scene-level datasets hinders the development of strong zero-shot generalizable 3D foundation models as opposed to their 2D counterparts. However, for comparison, we evaluate Swin3D [97], a 3D backbone that achieves leading performance in zero-shot perception tasks in multiple evaluation datasets compared to previous methods [91, 38, 109]. Swin3D is pretrained on Structured3D [113], a dataset 10 times larger than ScanNet [20]. In addition, we also evaluate the SAM model [43], an open-world instance segmentation model pretrained on the SA-1B [43] dataset, and the Uni3D model [116], which is an object-centric 3D foundation model pretrained on a mixture of datasets proposed by OpenShape [52]. The detailed results of the evaluation of these two models are provided in the supplementary material.
@@
Feature visualization.   Figure 3 visualizes the features of representative scenes extracted by the vision foundation models. To visualize a high-dimensional feature space with $C$ channels, we apply principal component analysis (PCA) to reduce the feature dimensions to three, normalize them to the range $[0,1]$, and interpret them as RGB color channels. We demonstrate several representative foundation models’ feature visualization, which reveals many intuitive findings. The image models, DINOv2 and LSeg, demonstrate strong semantic understanding, with LSeg exhibiting clearer discrimination due to its pixel-level language semantic guidance. The diffusion-based models, SD and SVD, in addition to their semantic modeling, excel at preserving the local geometry and texture of the scenes because of the generation-guided pretraining. The video models, SVD and V-JEPA, showcase a unique ability to identify different instances of the same semantic concepts, such as the two trees in the first scene and the chairs in both scenes. The 3D model, Swin3D, also exhibits strong semantic understanding. However, due to limited training data and domain shift, its quality is not on par with the image foundation models, despite being pretrained on perfect semantic annotations.
3.2 Vision-Language Reasoning
The vision-language reasoning task requires a model to engage in dialogues or answer questions about global understanding and local concepts related to a given complex 3D indoor scene. Following existing methods [36, 112], we formulate this as a visual-question answering (VQA) task using large language models (LLMs) as the backbone – given a 3D scene from multi-view images and point clouds, and a user-prompt question, the LLMs are asked to generate the answer to the question in an auto-regressive way. This task encompasses universal language-guided reasoning of the complex indoor scene, ranging from global layout to local details.
Datasets and optimization.   We evaluate the performance on two challenging indoor 3D VQA datasets: ScanQA [5] and SQA3D [55]. Following the evaluation methodology of [36, 55, 5, 58], we report the metrics BLEU [64], ROUGE [49], METEOR [7], and CIDEr [85]. We finetune a Q-Former module [48] to align features from different encoders to the LLM input space. More dataset and optimization details are provided in the supplementary material.
Evaluation results.   Table 2 and Figure 4 present the results of our evaluation. We observe that image and video encoders generally outperform the 3D point encoder, with DINOv2 achieving the best performance, followed closely by V-JEPA and SVD. Interestingly, we find that for LSeg and CLIP, which are pretrained by language guidance, their advantage in language alignment does not translate into superior performance on the LLM-guided VQA task. This finding challenges the common practice of using language-pretrained VFMs [68, 47, 48, 46] as default encoders for LLM-based vision-language reasoning tasks. Instead, it suggests the importance of considering a wider range of encoders, such as DINOv2 and V-JEPA, to support such tasks.